# Sorting the Babble in Babel: Assessing the Performance of Language Detection Algorithms on the OpenAlex Database

### v2

@@ 1 Introduction
At the crossroads of linguistics and computer science research, monolingual automatic language detection is nowadays considered a solved task in the literature, with state-of-the-art models achieving above human-level performance in as much as 1,366 languages (Caswell et al.,, 2020). Integrated in various real-world applications, language detection algorithms can also greatly enhance the organization, accessibility, and leveraging of information stored in databases. More precisely, by automatically identifying and categorizing the language of text attributes, these algorithms facilitate the sorting and indexing of multilingual data, which makes it easier to manage large, world-wide collections. Used in conjunction with machine translation algorithms, they can also support effective cross-language searches by allowing users to find pertinent information regardless of the language in which either the query or the relevant documents were written. Last but not least, language detection can benefit both data and resource management by respectively improving cross-linguistic data consistency and providing insights based on language distributions within the database.
@@
Despite these advantages, the use of language detection algorithms in bibliographic databases remains as rare as scarcely documented. The main reason for this gap in implementation and research is both historical and infrastructural, as databases such as Scopus or the Web of Science have traditionally and quasi-exclusively focused on publications written in English (Ammon,, 2006; Archambault et al.,, 2005; Mongeon and Paul-Hus,, 2016; Demeter,, 2020; Vera-Baceta et al.,, 2019). This linguistic bias, while greatly reducing the need for multi- and cross-linguistic functionalities, also impacts the quality and scope of both research evaluation and research on research: being mostly blind to anything that eludes their monolingual lens, traditional bibliometric databases thus provide a fundamentally skewed view of the disciplinary and geographical landscape of the global scientific community (Beigel and Digiampietri,, 2022), which as a consequence distorts every initiative or investigation it informs and whose repercussions affect the neglected communities in various ways and degrees (Tennant,, 2020; Salatino and Ruiz,, 2021; Finardi,, 2022; Navarro et al.,, 2022).
@@
Recent open-source alternative to the main traditional bibliometric databases, OpenAlex (OpAl) stands out from its predecessors by offering a broader landscape of scholarly communication (Priem et al.,, 2022). Built on an open codebase and from a variety of bibliometric projects, services, databases, and repositories, it has been shown to outperform both the Web of Science and Scopus in terms of document and journal frequency (Alperin et al.,, 2024; Culbert et al.,, 2024; van Bellen et al.,, 2024; Jiao et al.,, 2023). Of particular relevance here, recent research has shown that OpAl also has a higher proportion of articles written in non-English languages than the two above-mentioned bibliographic databases (Céspedes et al.,, 2025; Vera-Baceta et al.,, 2019). In light of these considerations, not only does OpAl have the potential to offer a broader, multilingual, and as such more representative view of scholarly production at the global scale, but its stronger allophone content makes it also more susceptible to benefit from the functionalities and advantages language detection algorithms can offer to multilingual databases.
@@
Regarding that last point, previous studies have however shown that language detection algorithm performance in multilingual context can vary significantly depending on the languages involved or on the specific characteristics of the dataset. In that perspective, effective multilingual database management requires linguistic classification procedures that are not only tailored to the specific linguistic features and metadata characteristics of the dataset, but also based on algorithms that are robust to noise in the data attributes considered (Balazevic et al.,, 2016; Polasi and Sri Rama Krishna,, 2016; Sujaini and Negara,, 2024). In the specific case of OpenAlex, while a quality assessment of OpAl language metadata has already been carried out (Céspedes et al.,, 2025), no attempt has yet been made yet to optimize the quality of its language metadata, that is, to find the best linguistic classification procedure for the articles it contains and indexes, based on its current metadata content and configuration.
@@ 2.1 Sample Collection
The ground truth for this study is based on the annotation work carried out by Céspedes et al., (2025) in their assessment of language metadata quality in OpAl. For that paper, a multilingual sample of indexed articles was manually annotated following a two step process. First, two samples of indexed articles were randomly drawn from the database, a first one containing 50 articles for each of the 55 languages present in the database (the resulting sample included only $2701$ articles, as some languages had less than 50 indexed articles) and a second containing 300 documents for each of the 11 most frequent languages in the database. Then, each article sampled was manually annotated by not only inspecting its title and abstract, but also by accessing the full text of the corresponding document, identify its language, and then determine if that paper indeed corresponds to a scientific paper; entries whose documents couldn’t be accessed were ignored (see (Céspedes et al.,, 2025) for more information about these samples).
@@ 2.1.1 Corpus Building
As this figure shows, most articles in the database are written in English, as reported in Céspedes et al., (2025), a majority of them also happen to include both abstracts and journal names as metadata. Following English, Asian languages form the most frequent linguistic group, followed by Romance languages. As for the different metadata attributes considered here, no clear cross-linguistic pattern can be found: while all articles have titles as attributes, metadata availability as regards to abstracts or journal names varies a lot across languages. Articles that include both their abstract and the name of the publishing journal represent a minority of indexed articles: besides English, only Chinese and Indonesian have a majority of indexed articles that include both articles and journal names. As for the remaining languages, most article entries only have journal names alongside titles, and a majority of Italian articles do not have any other relevant information other than titles.
@@ 2.2 Algorithms
All language detection algorithms whose performance is assessed here have been selected based on two criteria: Python implementability and completeness. The first criterion is not in itself problematic, since most language detection algorithms are implemented in Python or can be used in Python via packages specifically developed to that end. As for the second criterion, an algorithm is said to be complete if it terminates with a solution when one exists (Pearl,, 1984); in the case of the present project, any algorithm said to be complete is thus one that is able to return a linguistic prediction for each document given as input. Although this criterion does not seem too restrictive at first sight, some algorithms have a relatively low language coverage and have not been designed or trained to detect all the languages present in the OpAl database. For this and other reasons, any algorithm that was unable to return a prediction for every document submitted was excluded from subsequent analyses, regardless of its overall performance. Based on these restrictions, the different language detection algorithms selected and evaluated for this are shown and described in Table 1.
@@
CLD2 Second installment of the Compact Language Detector developed by Google for its Chrome browser, CLD2 is built on a Naïve Bayesian classifier trained on a manually constructed corpus of chosen web pages for each language (Sites,, 2025).
CLD3 Third installment of the Compact Language Detector designed for Chrome. It is a multi-layered neural network built on word embeddings based on character unigram, bigram, and trigram proportions and able to detect 107 languages (Salcianu et al.,, 2025).
DetectLanguage Partly based on Google’s CLD2 Cloud translation API, this language detection web service allows its users to detect the language of a text either through a POST request or its API (Detect Language API,, 2025).
FastText Created by Facebook’s AI Research lab for efficient learning of word representations and sentence classification, fastText’s language identification model is trained on data from Wikipedia, Tatoeba, and SETimes (Joulin et al., 2016b, ; Joulin et al., 2016a, ; Bojanowski et al.,, 2016). Available in both uncompressed and compressed versions, only the uncompressed and more performant version is used here.
FastSpell FastSpell complements FastText with spell-checking functionalities from Hunspell. This algorithm detects languages by first trying to determine the language of a sentence with FastText, then makes extra checks with Hunspell to determine the language more precisely (Bañón et al.,, 2024).
Langdetect A Java-based language-detection library, Langdetect is built on a Naive Bayes classifier, trained on a mix of n-gram-based language-independent character corpora extracted from Wikipedia abstracts (Shuyo,, 2010)
LangID This algorithm uses a Naive Bayes classifier trained on relative frequencies of character n-grams extracted from JRCAcquis, ClueWeb 09, Wikipedia, Reuters RCV2, and Debian i18n (Lui and Baldwin,, 2011; Heafield et al.,, 2015; Lui and Baldwin,, 2012)
@@
For every algorithm implemented, different linguistic classification procedures were designed by using it on every metadata corpus generated, resulting in a total of $4*7=28$ different sample partitions ($11*7=77$ in the appendix).
@@ 2.3 Performance Assessment
Following this and in accordance with the procedure used in Céspedes et al., (2025), true positive counts, false positive counts, and false negative counts were compiled and grouped by by classification procedure and language. Then, in order to assess the performance of each language detection procedure for each different language category, three different but complementary performance measures were implemented and used.
@@ 2.4.1 Robustness Assessment
As results presented in Céspedes et al., (2025) show, the quality of OpAl’s language metadata is surprisingly good; however, it it not perfect, which means that the language frequency distributions returned by the database are off to an unknown but certain extent. Also but perhaps most importantly, OpAl is an incessantly growing and changing database, as new articles are indexed and metadata of already indexed articles gets updated. The current state of the database in terms of language proportions and types of metadata is therefore not only uncertain at the present time, but also bound to change in the more or less short term. In that perspective, algorithmic robustness, that is, the ability of an algorithm to maintain ”some desired system characteristics despite fluctuations in the behavior of its component parts or its environment” (Carlson and Doyle,, 2002, p.2539), is bound to have a crucial impact of the quality of any linguistic classification of the database. In other words, an evaluation of linguistic classification procedures for the OpenAlex database has to take into account the ability of each procedure to that maintain good overall performance despite uncertainties in and changes to the current state of the database.
In order to best quantify database uncertainty and algorithm robustness, probabilistic confusion matrices were generated based on the sample predictions collected from the different language detection procedures. Several studies have already demonstrated the usefulness of Bayesian uncertainty modeling in the context of algorithm performance evaluation (Goutte and Gaussier,, 2005; Brodersen et al.,, 2010; Benavoli et al.,, 2017; Caelen,, 2017; Tötsch and Hoffmann,, 2021; Chlaily et al.,, 2023). The probabilistic approach adopted here will consist in modeling uncertainty by treating precision, recall as well as the relative proportion of languages and metadata types in the database as random variables modeled using different conjugate distribution models. In conformity with previous probabilistic modelisations of confusion matrices and submatrices (Brodersen et al.,, 2010; Tötsch and Hoffmann,, 2021), precision and recall probabilities for each algorithm, language, and metadata configuration, were modeled using posterior beta-binomial distributions with uniform priors for the confusion submatrices of each distinct database subgroup $k$; true positive counts were used as alpha parameters for both measures, while false positives and false negatives were used as beta parameters for precision and recall distributions respectively.
@@
Using these different conjugate distributional models, overall performance scores for the OpAl database were estimated through the following simulation procedure: for each classification procedure, precision and recall scores are obtained by drawing a random number from a mixture of beta-binomial distributions build on the classification outcomes for each language category, mixture whose component weights are obtained through a single draw of the language proportion model. This random sampling process is repeated 100 000100000100\,000100 000 times, and scores for each measure and procedure are compiled into a single simulation dataset.
@@ 2.4.2 Data Imbalance
As is often the case in large databases, OpenAlex suffers from severe distribution skews, many of which are caused by the preponderance of English-written articles, which allegedly account for $75.1$% of all database entries (imbalance ratio of 3:1). Such an asymmetry, apart from the question of its representativeness of the real linguistic distribution of scholarly communication on a global scale, poses a serious problem with regard to the evaluation of the linguistic classification procedures implemented for this project. As is the case with imbalanced sets in general, while the general objective is to improve recall without hurting precision, the number of false negatives for the minority classes can hardly be reduced without at the same time increasing their false positive count (He and Ma,, 2013). Moveover, standard classification algorithms are often biased towards the majority English class, leading to a higher misclassification rate instances belonging to minority classes, in the present case all other languages (López et al.,, 2013; Liu et al.,, 2006; Huang and Irwin,, 2006). In the specific context of the OpenAlex database and as pointed out in Céspedes et al., (2025), classifying article language in the OpenAlex database thus involves a trade-off between precision performance for English and retrieval performance for all other languages: a lower precision score for English will result in a better performance overall across all metrics, at the expense however of an underestimation of all other languages. This is all the more troublesome as the minority classes, i.e. all non-English languages in the present project, are usually those that are the most crucial from a classification perspective (López et al.,, 2013). In the context of the current project, it could also be argued that these minority classes are actually the most interesting ones, as they are what motivates the linguistic classification in the first place.
@@
Regarding $\beta$ values, $\beta>1$ and $\beta<1$, allows $F_{\beta,\gamma}$ to respectively confer more or less importance to recall in relation to precision, with lower and upper bounds $\beta=0.5$ and $\beta=2$ respectively conferring to recall half and twice the importance of precision. As for $\gamma$, $\gamma=0$ excludes speed from $F_{\beta,\gamma}$, $\gamma=1$ confers equal weight to both precision and speed, while lower and upper bounds $\gamma=0.5$ and $\gamma=2$ makes speed half and twice as important as precision, in that order. Given this, $\gamma=0$, allows $F_{\beta,\gamma}$ becomes formally equivalent to $F_{\beta}$, while setting $\beta=1$ and $\gamma=1$ makes $F_{\beta,\gamma}$ return the harmonic mean of precision, recall and speed. (Sasaki,, 2007).
Overall simulated performance for each procedure is then obtained by calculating $F_{\beta},\gamma$ scores obtained by each procedure at each simulation step for a set of various $\beta$ values ranging from $0.5$ (precision is twice as important as recall) to $2$ (recall is twice as important as precision) as well as 4 different values of $gamma$: 0 (speed is ignored altogether), 0.5 (speed is twice as important as recall), 1 (precision and speed are equally important), 2 (precision is twice as important as speed). Following these aggregative computations, two two different performance measures are then computed: maximum a posteriori estimates, which correspond to the mode of the posterior density obtained for each procedure, and win ratios, which are computed by dividing the number of times each procedure gets the best aggregated score by the total number of simulation iterations.

### v3

@@ 1 Introduction
At the crossroads of linguistics and computer science research, monolingual automatic language detection is nowadays considered a solved task in the literature, with state-of-the-art models achieving above human-level performance in as much as 1,366 languages (caswell2020language). Integrated in various real-world applications, language detection algorithms can also greatly enhance the organization, accessibility, and leveraging of information stored in databases. More precisely, by automatically identifying and categorizing the language of text attributes, these algorithms facilitate the sorting and indexing of multilingual data, which makes it easier to manage large, world-wide collections. Used in conjunction with machine translation algorithms, they can also support effective cross-language searches by allowing users to find pertinent information regardless of the language in which either the query or the relevant documents were written. Last but not least, language detection can benefit both data and resource management by respectively improving cross-linguistic data consistency and providing insights based on language distributions within the database.
@@
Despite these advantages, the use of language detection algorithms in bibliographic databases remains as rare as scarcely documented. The main reason for this gap in implementation and research is both historical and infrastructural, as databases such as Scopus or the Web of Science have traditionally and quasi-exclusively focused on publications written in English (ammon2006language; archambault2005welcome; mongeon2016journal; demeter2020academic; vera2019web). This linguistic bias, while greatly reducing the need for multi- and cross-linguistic functionalities, also impacts the quality and scope of both research evaluation and research on research: being mostly blind to anything that eludes their monolingual lens, traditional bibliometric databases thus provide a fundamentally skewed view of the disciplinary and geographical landscape of the global scientific community (beigel2022batalla), which as a consequence distorts every initiative or investigation it informs and whose repercussions affect the neglected communities in various ways and degrees (tennant2020web; salatino2021fetichismo; finardi2022linguas; navarro2022rethinking).
@@
Recent open-source alternative to the main traditional bibliometric databases, OpenAlex (OpAl) stands out from its predecessors by offering a broader landscape of scholarly communication (priem2022openalex). Built on an open codebase and from a variety of bibliometric projects, services, databases, and repositories, it has been shown to outperform both the Web of Science and Scopus in terms of document and journal frequency (alperin2024analysis; culbert2024reference; van2024oligopoly; jiao2023exclusively). Of particular relevance here, recent research has shown that OpAl also has a higher proportion of articles written in non-English languages than the two above-mentioned bibliographic databases (cespedes2025evaluating; vera2019web). In light of these considerations, not only does OpAl have the potential to offer a broader, multilingual, and as such more representative view of scholarly production at the global scale, but its stronger allophone content makes it also more susceptible to benefit from the functionalities and advantages language detection algorithms can offer to multilingual databases.
@@
Regarding that last point, previous studies have however shown that language detection algorithm performance in multilingual context can vary significantly depending on the languages involved or on the specific characteristics of the dataset. In that perspective, effective multilingual database management requires linguistic classification procedures that are not only tailored to the specific linguistic features and metadata characteristics of the dataset, but also based on algorithms that are robust to noise in the data attributes considered (balazevic2016language; polasi2016combining; sujaini2024). In the specific case of OpenAlex, while a quality assessment of OpAl language metadata has already been carried out (cespedes2025evaluating), no attempt has yet been made yet to optimize the quality of its language metadata, that is, to find the best linguistic classification procedure for the articles it contains and indexes, based on its current metadata content and configuration.
@@ 2.1 Sample Collection
The ground truth for this study is based on the annotation work carried out by cespedes2025evaluating in their assessment of language metadata quality in OpAl. For that paper, a multilingual sample of indexed articles was manually annotated following a two step process. First, two samples of indexed articles were randomly drawn from the database, a first one containing 50 articles for each of the 55 languages present in the database (the resulting sample included only $2701$ articles, as some languages had less than 50 indexed articles) and a second containing 300 documents for each of the 11 most frequent languages in the database. Then, each article sampled was manually annotated by not only inspecting its title and abstract, but also by accessing the full text of the corresponding document, identify its language, and then determine if that paper indeed corresponds to a scientific paper; entries whose documents couldn’t be accessed were ignored (see (cespedes2025evaluating) for more information about these samples).
@@ 2.1.1 Corpus Building
As this figure shows, most articles in the database are written in English, as reported in cespedes2025evaluating, a majority of them also happen to include both abstracts and journal names as metadata. Following English, Asian languages form the most frequent linguistic group, followed by Romance languages. As for the different metadata attributes considered here, no clear cross-linguistic pattern can be found: while all articles have titles as attributes, metadata availability as regards to abstracts or journal names varies a lot across languages. Articles that include both their abstract and the name of the publishing journal represent a minority of indexed articles: besides English, only Chinese and Indonesian have a majority of indexed articles that include both articles and journal names. As for the remaining languages, most article entries only have journal names alongside titles, and a majority of Italian articles do not have any other relevant information other than titles.
@@ 2.2 Algorithms
All language detection algorithms whose performance is assessed here have been selected based on two criteria: Python implementability and completeness. The first criterion is not in itself problematic, since most language detection algorithms are implemented in Python or can be used in Python via packages specifically developed to that end. As for the second criterion, an algorithm is said to be complete if it terminates with a solution when one exists (pearl1984heuristics); in the case of the present project, any algorithm said to be complete is thus one that is able to return a linguistic prediction for each document given as input. Although this criterion does not seem too restrictive at first sight, some algorithms have a relatively low language coverage and have not been designed or trained to detect all the languages present in the OpAl database. For this and other reasons, any algorithm that was unable to return a prediction for every document submitted was excluded from subsequent analyses, regardless of its overall performance. Based on these restrictions, the different language detection algorithms selected and evaluated for this are shown and described in Table 1.
@@
CLD2 Second installment of the Compact Language Detector developed by Google for its Chrome browser, CLD2 is built on a Naïve Bayesian classifier trained on a manually constructed corpus of chosen web pages for each language (sites2025compact).
CLD3 Third installment of the Compact Language Detector designed for Chrome. It is a multi-layered neural network built on word embeddings based on character unigram, bigram, and trigram proportions and able to detect 107 languages (salcianu2025).
DetectLanguage Partly based on Google’s CLD2 Cloud translation API, this language detection web service allows its users to detect the language of a text either through a POST request or its API (detectlanguage2025).
FastText Created by Facebook’s AI Research lab for efficient learning of word representations and sentence classification, fastText’s language identification model is trained on data from Wikipedia, Tatoeba, and SETimes (joulin2016bag; joulin2016fasttext; bojanowski2016enriching). Available in both uncompressed and compressed versions, only the uncompressed and more performant version is used here.
FastSpell FastSpell complements FastText with spell-checking functionalities from Hunspell. This algorithm detects languages by first trying to determine the language of a sentence with FastText, then makes extra checks with Hunspell to determine the language more precisely (banon2024fastspell).
Langdetect A Java-based language-detection library, Langdetect is built on a Naive Bayes classifier, trained on a mix of n-gram-based language-independent character corpora extracted from Wikipedia abstracts (nakatani2010langdetect)
LangID This algorithm uses a Naive Bayes classifier trained on relative frequencies of character n-grams extracted from JRCAcquis, ClueWeb 09, Wikipedia, Reuters RCV2, and Debian i18n (lui2011cross; heafield2015language; lui2012langid)
@@
For every algorithm implemented, different linguistic classification procedures were designed by using it on every metadata corpus generated, resulting in a total of 47=284*7=28 different sample partitions (117=7711*7=77 in the appendix).
@@ 2.3 Performance Assessment
Following this and in accordance with the procedure used in cespedes2025evaluating, true positive counts, false positive counts, and false negative counts were compiled and grouped by by classification procedure and language. Then, in order to assess the performance of each language detection procedure for each different language category, three different but complementary performance measures were implemented and used.
@@ 2.4.1 Robustness Assessment
As results presented in cespedes2025evaluating show, the quality of OpAl’s language metadata is surprisingly good; however, it it not perfect, which means that the language frequency distributions returned by the database are off to an unknown but certain extent. Also but perhaps most importantly, OpAl is an incessantly growing and changing database, as new articles are indexed and metadata of already indexed articles gets updated. The current state of the database in terms of language proportions and types of metadata is therefore not only uncertain at the present time, but also bound to change in the more or less short term. In that perspective, algorithmic robustness, that is, the ability of an algorithm to maintain ”some desired system characteristics despite fluctuations in the behavior of its component parts or its environment” (carlson2002complexity, p.2539), is bound to have a crucial impact of the quality of any linguistic classification of the database. In other words, an evaluation of linguistic classification procedures for the OpenAlex database has to take into account the ability of each procedure to that maintain good overall performance despite uncertainties in and changes to the current state of the database.
In order to best quantify database uncertainty and algorithm robustness, probabilistic confusion matrices were generated based on the sample predictions collected from the different language detection procedures. Several studies have already demonstrated the usefulness of Bayesian uncertainty modeling in the context of algorithm performance evaluation (goutte2005probabilistic; brodersen2010balanced; benavoli2017time; caelen2017bayesian; totsch2021classifier; chlaily2023measures). The probabilistic approach adopted here will consist in modeling uncertainty by treating precision, recall as well as the relative proportion of languages and metadata types in the database as random variables modeled using different conjugate distribution models. In conformity with previous probabilistic modelisations of confusion matrices and submatrices (brodersen2010balanced; totsch2021classifier), precision and recall probabilities for each algorithm, language, and metadata configuration, were modeled using posterior beta-binomial distributions with uniform priors for the confusion submatrices of each distinct database subgroup $k$; true positive counts were used as alpha parameters for both measures, while false positives and false negatives were used as beta parameters for precision and recall distributions respectively.
@@
Using these different conjugate distributional models, overall performance scores for the OpAl database were estimated through the following simulation procedure: for each classification procedure, precision and recall scores are obtained by drawing a random number from a mixture of beta-binomial distributions build on the classification outcomes for each language category, mixture whose component weights are obtained through a single draw of the language proportion model. This random sampling process is repeated 100 000100\,000 times, and scores for each measure and procedure are compiled into a single simulation dataset.
@@ 2.4.2 Data Imbalance
As is often the case in large databases, OpenAlex suffers from severe distribution skews, many of which are caused by the preponderance of English-written articles, which allegedly account for 75.175.1% of all database entries (imbalance ratio of 3:1). Such an asymmetry, apart from the question of its representativeness of the real linguistic distribution of scholarly communication on a global scale, poses a serious problem with regard to the evaluation of the linguistic classification procedures implemented for this project. As is the case with imbalanced sets in general, while the general objective is to improve recall without hurting precision, the number of false negatives for the minority classes can hardly be reduced without at the same time increasing their false positive count (he2013imbalanced). Moveover, standard classification algorithms are often biased towards the majority English class, leading to a higher misclassification rate instances belonging to minority classes, in the present case all other languages (lopez2013insight; liu2006boosting; huang2006intelligent). In the specific context of the OpenAlex database and as pointed out in cespedes2025evaluating, classifying article language in the OpenAlex database thus involves a trade-off between precision performance for English and retrieval performance for all other languages: a lower precision score for English will result in a better performance overall across all metrics, at the expense however of an underestimation of all other languages. This is all the more troublesome as the minority classes, i.e. all non-English languages in the present project, are usually those that are the most crucial from a classification perspective (lopez2013insight). In the context of the current project, it could also be argued that these minority classes are actually the most interesting ones, as they are what motivates the linguistic classification in the first place.
@@
Regarding $\beta$ values, $\beta>1$ and $\beta<1$, allows $F_{\beta,\gamma}$ to respectively confer more or less importance to recall in relation to precision, with lower and upper bounds $\beta=0.5$ and $\beta=2$ respectively conferring to recall half and twice the importance of precision. As for $\gamma$, $\gamma=0$ excludes speed from $F_{\beta,\gamma}$, $\gamma=1$ confers equal weight to both precision and speed, while lower and upper bounds $\gamma=0.5$ and $\gamma=2$ makes speed half and twice as important as precision, in that order. Given this, $\gamma=0$, allows $F_{\beta,\gamma}$ becomes formally equivalent to $F_{\beta}$, while setting $\beta=1$ and $\gamma=1$ makes $F_{\beta,\gamma}$ return the harmonic mean of precision, recall and speed. (sasaki2007truth).
Overall simulated performance for each procedure is then obtained by calculating $F_{\beta},\gamma$ scores obtained by each procedure at each simulation step for a set of various $\beta$ values ranging from $0.5$ (precision is twice as important as recall) to $2$ (recall is twice as important as precision) as well as 4 different values of gammagamma: 0 (speed is ignored altogether), 0.5 (speed is twice as important as recall), 1 (precision and speed are equally important), 2 (precision is twice as important as speed). Following these aggregative computations, two two different performance measures are then computed: maximum a posteriori estimates, which correspond to the mode of the posterior density obtained for each procedure, and win ratios, which are computed by dividing the number of times each procedure gets the best aggregated score by the total number of simulation iterations.